2x+4y=24
-15x+10y=-20
x=? y=?

Answer by checkley77(12844)

2x+4y=24 multiply this equation by -2.5 & add.
-15x+10y=-20
-5x-10y=-60
---------------------
-20x=-80
x=-80/-20
x=4 answer.
2*4+4y=24
8+4y=24
4y=24-8
4y=16
y=16/4
y=4 answer.